OpenStudy (chihiroasleaf):

Can anyone explain to me about solving partial differential equation with characteristic method? or any reference to study... I really have no idea about this...

OpenStudy (bittuaryan):

enclosed is a pdf link, may it helpful, http://www.math.ucsb.edu/~grigoryan/124A.pdf

OpenStudy (chihiroasleaf):

thank you... :)

OpenStudy (bittuaryan):

not mention.....please read at first & then tell us.

OpenStudy (chihiroasleaf):

I don't really understand when the right side of the differential equation is not zero..., can you explain it?

OpenStudy (bittuaryan):

If it is not '0' in right side, you make it by transfer that value in left side.

OpenStudy (chihiroasleaf):

so.., if I have PDE \( \Large \frac{\partial u }{\partial t} + x \frac{\partial u}{\partial x} = 1\) should I move the '1' to the left side? so far I get the characteristic equation \( \frac{dx}{dt} = x \) and \( \frac{du}{dt} = 1 \) and end up with getting \( x = xt+ x_0 \) and \(u = t + u_0\) and have no idea what to do next... is it correct?